(Present value​) What is the present value of the following future​ amounts?

a.  ​$800 to be received 8 years from now discounted back to the present at 11 percent.

b.  ​$200 to be received 7 years from now discounted back to the present at 7 percent.

c.  ​$1150 to be received 13 years from now discounted back to the present at 3 percent.

d.  ​$1250 to be received 4 years from now discounted back to the present at 18 percent.
Answer #1

Calculation of present value:

Present value= Future value*PVF@rate for number of years

a) Calculation of present value:

Present value= 800*PVF@11% for 8 years= 800*0.4339= $347.12

b) calculation of present value:

Present value= 200*PVF @7% for 7 years

Present value= 200*0.6227= $124.54

c) Calculation of present value:

Present value= 1150*PVF @3% for 13 years

Present value= 1150*0.6810= $783.15

d) Calculation of present value:

Present value= 1250*PVF@18% for 4 years

Present value= 1250*0.7350= $918.75